OpenStudy (anonymous):

radical x + radical 2x =1

OpenStudy (anonymous):

\[=\sqrt{x}+\sqrt{2x}=1\]

OpenStudy (anonymous):

\[\sqrt{x}*(1+\sqrt{2})=1\]

OpenStudy (anonymous):

\[\sqrt{x}=1/(1+\sqrt{2})\] square both side - you get answer for x

OpenStudy (anonymous):

\[x=1/(1+\sqrt{2})^{2}\]

OpenStudy (anonymous):

so you can't combine \[\sqrt{x}\] and \[\sqrt{2x}\]